# Trustworthy Reasoning: Evaluating and Enhancing Factual Accuracy in LLM Intermediate Thought Processes

### v1

@@ Abstract
We present RELIANCE (Reasoning Evaluation with Logical Integrity and Accuracy for Confidence Enhancement), a novel framework addressing a critical vulnerability in Large Language Models (LLMs): the prevalence of factual inaccuracies within intermediate reasoning steps despite correct final answers. This phenomenon poses substantial risks in high-stakes domains including healthcare, legal analysis, and scientific research, where erroneous yet confidently presented reasoning can mislead users into dangerous decisions. Our framework integrates three core components: (1) a specialized fact-checking classifier trained on counterfactually augmented data to detect subtle factual inconsistencies within reasoning chains; (2) a Group Relative Policy Optimization (GRPO) reinforcement learning approach that balances factuality, coherence, and structural correctness through multi-dimensional rewards; and (3) a mechanistic interpretability module examining how factuality improvements manifest in model activations during reasoning processes. Extensive evaluation across ten state-of-the-art models reveals concerning patterns: even leading models like Claude-3.7 and GPT-o1 demonstrate reasoning factual accuracy of only 81.93% and 82.57% respectively. RELIANCE significantly enhances factual robustness (up to 49.90% improvement) while maintaining or improving performance on challenging benchmarks including Math-500, AIME-2024, and GPQA. Furthermore, our activation-level analysis provides actionable insights into how factual enhancements reshape reasoning trajectories within model architectures, establishing foundations for future training methodologies that explicitly target factual robustness through activation-guided optimization.
@@ 5.2 Factuality Enhancement Evaluation (RQ2)
Factual Reliability: While our RELIANCE framework substantially enhances factual reasoning steps across models, a natural question arises: does this reasoning steps factuality improvement compromise the quality of the final reasoning output? To address this, we conducted a comprehensive evaluation across several benchmark datasets. Our analysis reveals that RELIANCE not only maintains task performance but can also yield modest improvements, while significantly reducing factual errors. As shown in Figure 2, the DeepSeek-R1-Distill-Qwen-1.5B model augmented with RELIANCE achieves slightly higher accuracy on Math-500 (83.16% vs. 82.42%) and AIME-2024 (26.67% vs. 26.53%) compared to the baseline. Performance remains largely consistent on GPQA (33.22% vs. 33.84%) and LiveCodeBench (15.51% vs. 16.13%). Finding #3: The RELIANCE framework significantly enhances factual reasoning accuracy and consistency across various open-source white-box language models, with the most dramatic improvements observed in smaller models. Importantly, RELIANCE not only boosts factual reasoning but also maintains or slightly improves downstream performance on real-world benchmarks such as Math-500, AIME-2024, GPQA, and LiveCodeBench. These results demonstrate that RELIANCE enhances factual reliability without compromising the utility of models across diverse reasoning tasks.

### v2

@@ Abstract
We present a novel framework addressing a critical vulnerability in Large Language Models (LLMs): the prevalence of factual inaccuracies within intermediate reasoning steps despite correct final answers. This phenomenon poses substantial risks in high-stakes domains including healthcare, legal analysis, and scientific research, where erroneous yet confidently presented reasoning can mislead users into dangerous decisions. Our framework integrates three core components: (1) a specialized fact-checking classifier trained on counterfactually augmented data to detect subtle factual inconsistencies within reasoning chains; (2) an enhanced Group Relative Policy Optimization (GRPO) reinforcement learning approach that balances factuality, coherence, and structural correctness through multi-dimensional rewards; and (3) a mechanistic interpretability method examining how factuality improvements manifest in model activations during reasoning processes. Extensive evaluation across multi state-of-the-art models reveals concerning patterns: even leading models like Claude-3.7 and GPT-o1 demonstrate reasoning factual accuracy of only 81.93% and 82.57% respectively. Our approach significantly enhances factual robustness (up to 49.90% improvement) while maintaining or improving performance on challenging benchmarks including Math-500, AIME-2024, and GPQA. Furthermore, our neural activation-level analysis provides actionable insights into how factual enhancements reshape reasoning trajectories within model architectures, establishing foundations for future training methodologies that explicitly target factual robustness through activation-guided optimization.
@@ 5.2 Factuality Enhancement Evaluation (RQ2)
Factual Reliability: While our RELIANCE framework substantially enhances factual reasoning steps across models, a natural question arises: does this reasoning steps factuality improvement compromise the quality of the final reasoning output? To address this, we conducted a comprehensive evaluation across several benchmark datasets. Our analysis reveals that RELIANCE not only maintains task performance but can also yield modest improvements, while significantly reducing factual errors. As shown in Figure 2, the DeepSeek-R1-Distill-Qwen-1.5B model augmented with RELIANCE achieves slightly higher accuracy on Math-500 (83.16% vs. 82.42%) and AIME-2024 (26.67% vs. 26.53%) compared to the baseline. Performance remains largely consistent on GPQA (34.22% vs. 33.84%) and LiveCodeBench (17.51% vs. 16.13%). Finding #3: The RELIANCE framework significantly enhances factual reasoning accuracy and consistency across various open-source white-box language models, with the most dramatic improvements observed in smaller models. Importantly, RELIANCE not only boosts factual reasoning but also maintains or slightly improves downstream performance on real-world benchmarks such as Math-500, AIME-2024, GPQA, and LiveCodeBench. These results demonstrate that RELIANCE enhances factual reliability without compromising the utility of models across diverse reasoning tasks.